LOG-IN
Displaying report 1-1 of 1.
MIF (General)
Print this report.
kazuya.yokogawa - 21:55, Thursday 09 August 2018 (5772)Get code to link to this report
Installation of ALS system

[Izumi, Sugimoto, Yokogawa]

Inspection of delivered optical fibers

We inspected delivered optical fibers(Fig1).

First of all, we aligned using 5m fiber for test. Then the transmittance was about 67%.

On the other hand, the transmittance of delivered optical fibers(58mx2 and 63mx2) were about 30% and 40% respectively. (Izumi-san will comment on details.)

Although these values are not good, the transmission reach 50% if the coupling loss is deducted. So, optical fibers laying work was started.

The Prometheus laser problem

The Prometheus laser suddenly shut down when I tried to rise the laser current from 1.0A to 2.2A. (The nominal current is 2.2A.)

After a while, I found the laser diode temperature Guard lamps was lighting. (Fig2)

According to the manual, these lamps shine if the laser diode temperature is lower than the room temperature. So, we changed the temperature of diode A from 24deg to 25deg, and we were able to rise the laser current to 2.0A.

However, this problem has not yet been solved, and we will continue to look at the situation tomorrow.

Images attached to this report
5772_20180809145314_img1599.jpg 5772_20180809145333_img1606.jpg
Comments related to this report
ryousuke.sugimoto - 00:42, Friday 10 August 2018 (5773)
I have a comment about PSL room's optical table.
Today, I found GND of power supply is conducting to optical table.
kiwamu.izumi - 06:09, Friday 10 August 2018 (5774)

A summary of the fiber transmittance measurements

Yokogawa, Sugimoto, Izumi

The below is a summary table of the fiber transmittance measurements performed today (the 9th) before the fibers were layed. The measurement was done by using the setup for the X arm green beam.

  Measured output power Estimated core transmittance*
CCH-948-63 A (63 m duplex fiber, A) 14 mW (36%)  60%
CCH-948-63 B (63 m duplex fiber, B) 12 mW (31%)  51%
CCH-948-58 (58 m duplex fiber, A) 16 mW (41%)  68%
CCH-948-58 (58 m duplex fiber, B) 16 mW (41%)  68%

* core tranmissttance = (measured output power) / (incident power) / (coupling efficiency). The input power was set to 39 mW and the coupliing efficiency was assumed to be 60%.

The fibers themself are made of Nufern PM-S405-XP whose core attenuation are nominally 30 dB/km @630 nm and @488 nm. From this value we were expecting the transmittance to be at least 65 % ( = 10^{-30 dB/1000 m x 58 m / 1000 m / 10}) in power for the 63 m fibres while 67% for the 58 m fibers. While the 58 m fibers barely meet the catalog spefication, the 63 m fibers do not match the catalog specification.

One thing we thought after the measurements was that we could have tried to adjust the input polarization. So, there may be some small room to improve the transmittance.

kiwamu.izumi - 09:21, Friday 10 August 2018 (5775)

A brief update on the ALS laser issue

This is a good news.

In this morning, we confirmed that the Prometheus (2143A, the one had the issue of suddenly shutting itself down when the current is high) stayed on overnight.

To clarify the situation, here is a list of things that happened or we did in the chronological order.

  • Aug.8th: We brought 2143A from the POP table to the PSL table.
  • Aug.9th, afternoon: The issue was discovered.
  • Aug.9th, late evening: The temperature guard indicators were found to be on for both the A and B diodes.
  • Aug.9th, late evening: Following the instruction written in the manual, we cranked up the temperature set point of diode A from the nominal of 24.05 deg to 25.00 deg. After this, we were able to raise the current as high 2.0 A and left the laser on.
  • Aug.10th (today): 2143A was found to have stayed on overnight without shutting itself down.
  • Aug.10th (today): The temperature setpoint of diode B was also cranked up from the nominal of 26.01 deg to 26.5 deg. So far so good.

The latest settings:

  • The temperature setpoint for diode A = 25.0 deg (c.f., nominal is 24.05 deg)
  • The temperature setpoint for diode B = 26.5 deg (c.f., nominal is 26.01 deg)
  • The diode injection current = 2.2 A, the nominal value.
  • Any other parameters were set exactly to the values listed in the data-sheet/test-report.
kiwamu.izumi - 10:16, Friday 10 August 2018 (5776)

FYI:

The test reports from the manufacturer are now avialbe in DCC at

https://gwdoc.icrr.u-tokyo.ac.jp/cgi-bin/private/DocDB/ShowDocument?docid=8692

in which one can check the nominal settings for the parameters including the diode temperature and such. The DCC page is password-protected.